# Frustration-free free fermions

### v1

@@ Appendix D Interacting bound of the neutral excitation gap
In our finite-size analysis, we take the linear size to be $L=3n+1$ or $L=3n+2$ with integer $n\geq 2$ such that we avoid sampling the K and K’ points. We parameterize $\bm{q}=\bm{q}_{\rm K}+\delta\bm{q}$ with $\delta\bm{q}=\delta q(\cos\theta,\sin\theta)$. For any fixed $\theta$, one can verify that $e_{\bm{q}_{\rm K}+\delta\bm{q}}$ increases monotonically with $70<\delta q\leq 1/$ 7. In the finite-size sampling of the BZ, there must be at least one sampled momentum within a distance of $1/L$ from the K point. We can therefore bound $\min_{\bm{q}}e_{\bm{q}}\leq\max_{\theta}e_{\bm{q}_{K}+(\cos\theta,\sin\theta)/L}$. In the vicinity of the K point, the single-particle dispersion is maximized at $\theta=5\pi/4$, and we may evaluate explicitly to find
@@
Appendix E Data analysis
Figure 4: Fitting of numerical data. Red crosses indicated numerical data and dashed black lines are the linear fits. (a) The linear fit of $t/E^{\rm SMA}(\delta\bm{k})$ against $\log L$ gives a slope of $1.04(2)$ and an intercept of $-1.179(7)$. (b) The linear fit of the $\log$ of the minimum particle-hole excitation energy at $\bm{k}=\bm{0}$ against $\log L$ gives a slope of $-1.99(2)$ and an intercept of $2.84(8)$.

### v2

@@ Appendix B Details of the honeycomb model
In our finite-size analysis, we take the linear size to be $L=3n+1$ or $L=3n+2$ with integer $n\geq 2$ such that we avoid sampling the K and K’ points. We parameterize $\bm{q}=\bm{q}_{\rm K}+\delta\bm{q}$ with $\delta\bm{q}=\delta q(\cos\theta,\sin\theta)$. For any fixed $\theta$, one can verify that $e_{\bm{q}_{\rm K}+\delta\bm{q}}$ increases monotonically with $0<\delta q\leq 1/7$. In the finite-size sampling of the BZ, there must be at least one sampled momentum within a distance of $1/L$ from the K point. We can therefore bound $\min_{\bm{q}}e_{\bm{q}}\leq\max_{\theta}e_{\bm{q}_{K}+(\cos\theta,\sin\theta)/L}$. In the vicinity of the K point, the single-particle dispersion is maximized at $\theta=5\pi/4$, and we may evaluate explicitly to find
@@
Figure 3: Fitting of numerical data. Red crosses indicated numerical data and dashed black lines are the linear fits. (a) The linear fit of $t/E^{\rm SMA}(\delta\bm{k})$ against $\log L$ gives a slope of $1.04(2)$ and an intercept of 1.179(7)-1.179(7). (b) The linear fit of the $\log$ of the minimum particle-hole excitation energy at $\bm{k}=\bm{0}$ against $\log L$ gives a slope of 1.99(2)-1.99(2) and an intercept of $2.84(8)$.